Today we spent time dealing with bolo electronics -- an unexpected reboot of the crates overnight meant that the overnight tuning didn't work, and later we had some other board issues that we've decided to debug more fully before proceeding further. That took up most of the day for myself, Jeff, and Kevin. In addition, Jeff has spent a lot of time refining his data analysis of our polarized Ebert-Fastie data.
The gondola team has made great progress today, mounting both mirrors and the primary mirror protection rollbar along with a few other ancillary pieces of gondola hardware (gyro box, star camera mounts, elevation actuator, etc). Tomorrow they plan on setting the inner frame atop the outer frame, joining the two major pieces of the gondola together again.
In other news, I've decided to move all BLAST-Pol pictures into a separate web album to avoid confusion with EBEX hardware. I will continue to update it but I probably won't post the link here again unless there's something particularly notable. You can see that album here:
Today's (EBEX) pictures: